OpenStudy (anonymous):

find the length of the missing side a.35 b.49 c.7 d.2

OpenStudy (anonymous):

@csscgirl

OpenStudy (***[isuru]***):

use the Pythagoras theorem...

OpenStudy (***[isuru]***):



OpenStudy (anonymous):

7 ?

OpenStudy (***[isuru]***):

yes!

OpenStudy (anonymous):

what about this question how do i do it ? find the length of the missing side leave answer in simplest radical form a. Sqrt 17 b. 133m c. sqrt 133m d. sqrt 71 m

OpenStudy (***[isuru]***):

\[C^{2} =A^{2} + B^{2}\]

OpenStudy (***[isuru]***):

\[C^{2} = 8^{2} + 7^{2}\]

OpenStudy (***[isuru]***):

\[C = \sqrt{8^{2} + 7^{2} }\]

OpenStudy (***[isuru]***):

wt will u get if u solve this ?

OpenStudy (anonymous):

i got 133

OpenStudy (***[isuru]***):

C^2 = 133 then C = ?

OpenStudy (anonymous):

its either b or c

OpenStudy (***[isuru]***):

\[C^{2} = 133\] \[C = \sqrt {133}\] got it :) ?

OpenStudy (anonymous):

yes thanks

OpenStudy (***[isuru]***):

u r welcome!